A certain charge Q is devided in two parts q & Q-q .for max force between them.find ratio q/Q

Grade:12

1 Answers

siddharth gupta
9 years ago
using  A.M-G.M inequality: for maximum force between two charges q(Q-q) should be maximum accordingly q=Q-q ie 2q=Q or q/Q=1/2.
WITH REGARDS
SIDDHARTH GUPTA